# A Diffusion-based generative equalizer for music restoration

### v1

@@ 5 Discussion: Restoring historical voices
One of the key contributions from our research lies in our efforts to restore early-20th-century recordings of classical opera singers. These recordings are often so much degraded that some of the unique vocal characteristics of the singer are no longer preserved. The role of BABE-2 is to reconstruct, or hallucinate, the missing pieces, based on statistical patterns derived from the training data. This underscores the critical importance of selecting appropriate training data, as it significantly impacts the accuracy and authenticity of the audio restoration, ensuring the reconstructed voices closely mirror the original timbres and nuances.
@@ 5.2 Qualitative Analysis
The white boxes in Fig. 4(c, d) show some of the consequences of the source separation preprocessing. The bleeding from the instrumental track is present in the restored recording, especially in the low-frequency region. Nevertheless, we also notice that a side-effect of BABE-2 is that it additionally suppresses some of the bleeding between the harmonics, producing an even cleaner vocal separation. This behavior is expected, as these artifacts are not included in the training data.
@@ 7 Acknowledgments
We acknowledge the computational resources provided by the Aalto Science-IT project. We thank Bryn Louise for proofreading parts of this paper and professor Anne Kauppala for perspectives on historical recordings. Maija Turunen’s research is funded by the Kone Foundation.
@@ A.1 Training
Table 2: Hyperparameters chosen in our experiments
@@ A.3 Inference
Table 3: Hyperparameters chosen in our experiments
@@ A.4 Filter Parameterization
The proposed filter parameterization is detailed in Sec. 3.1. For completeness, we report its mathematical formulation, which is as follows:
@@ A.5 Block-Autoregressive Inference
This adaptation can be equivalently applied to the filter reconstruction cost function in Eq. (LABEL:eq:filter_rec), allowing us to estimate the filter parameters with the updated objective. Note however that, if the degradation is considered to be time-invariant, there is no need to re-estimate the filter parameters $\phi$ among different frames, as only once is needed. Nevertheless, we use the block-autoregressive inference in our experiments for evaluation purposes.

### v2

@@ 5 Discussion: Restoring historical voices
One of the key contributions from our research lies in our efforts to restore early-20th-century recordings of classical opera singers. These recordings are often so much degraded that some of the unique vocal characteristics of the singer are no longer preserved. The role of BABE-2 is to reconstruct, or hallucinate, the missing pieces, based on statistical patterns derived from the training data. As a result, the final output can be considered a hybrid containing features from both the original recording and the training data. This underscores the critical importance of selecting appropriate training data, as it significantly impacts the accuracy and authenticity of the audio restoration, ensuring the reconstructed voices closely mirror the original timbres and nuances.
@@ 5.2 Qualitative Analysis
The white boxes in Fig. 4(c, d) show some of the consequences of the source separation preprocessing. The bleeding from the instrumental track is present in the restored recording, especially in the low-frequency region. The lack of robustness of the source separation model is a limitation that should be approached in future work. Nevertheless, we also notice that a side-effect of BABE-2 is that it additionally suppresses some of the bleeding between the harmonics, producing an even cleaner vocal separation. This behavior is expected, as these artifacts are not included in the training data.
@@ 7 Acknowledgments
We acknowledge the computational resources provided by the Aalto Science-IT project. We thank Bryn Louise for proofreading parts of this paper.
@@ A.1 Training
Table 2: Training hyperparameters chosen in our experiments
@@ A.3 Inference
Table 3: Inference hyperparameters chosen in our experiments
@@ A.4 Filter Parameterization
The proposed filter parameterization is introduces in Sec. 3.1. The piece-wise linear transfer function can be formalized as follows:
@@ A.5 Block-Autoregressive Inference
This adaptation can be equivalently applied to the filter reconstruction cost function in Eq. (13), allowing us to estimate the filter parameters with the updated objective. Note however that, if the degradation is considered to be time-invariant, there is no need to re-estimate the filter parameters $\phi$ among different frames, as only once is needed. Nevertheless, we use the block-autoregressive inference in our experiments for evaluation purposes.